OpenStudy (anonymous):

Reaction of the anti-cancer drug cisplatin, Pt(NH3)2Cl2, with water is described by the equation Pt(NH3)2Cl2(aq)+H2O(l)→ Pt(NH3)2(H2O)Cl+(aq)+Cl−(aq) The rate of this reaction increases by a factor of 15 on raising the temperature from 25 ∘C to 50 ∘C. what is the activation energy?

OpenStudy (aaronq):

use the arrhenius equation: \(ln(\dfrac{k_1}{k_2})=\dfrac{E_a}{R}(\dfrac{1}{T_1}-\dfrac{1}{T_2})\)

OpenStudy (anonymous):

what is the rate constant in this problem though? 15?

OpenStudy (aaronq):

it depends, if k1=1, then k2=15, or you can writ them the other way around.

OpenStudy (aaronq):

just make sure you're consistent with the temperatures.

OpenStudy (anonymous):

if K1=1 and K2=15, then t1=232 and t2=298?

OpenStudy (aaronq):

yeah, thats right. Make sure you write k in lower case, because capital K is the equilibrium constant.